# IMUFace: Real-Time, Low-Power, Continuous 3D Facial Reconstruction Through Earphones

### v1

@@ V-D4 Latency and Power Consumption
In addition, we used a power monitoring device to measure the power consumption of IMUFace. Specifically, when the system is idle, that is, the microcontroller unit (MCU) is in sleep mode and does not transmit data via Bluetooth, the average power consumption of IMUFace is 50 mW (13.5 mA @ 3.70 V). When IMUFace actively collects IMU data and transmits it via Bluetooth, the total system power consumption is 58 mW (15.7 mA @ 3.70 V). This indicates that with a 100 mAh lithium polymer battery, IMUFace can continuously record data for up to 6.5 hours, meeting the needs of most applications. We examined the power consumption of recent studies on 3D facial reconstruction and summarized the overall power consumption of the system, including the use of sensing and communication power, in Table II. Our analysis shows that IMUFace consumes the least power among all sensing technologies evaluated, with energy consumption 58.0% lower than that of the next most efficient technology, Bioface-3D [19].
TABLE II: Comparison of System Power Consumption among Related Existing Studies
@@
NeckFace [12] Cameras 4 W
C-Face [22] Cameras >4 W
BioFace-3D [19] Biosensors 138 mW
EarIO [20] Acoustics 153.7 mW
EyeEcho [15] Acoustics 167 mW
@@
VI Discussion
Our study aims to use conventional motion sensors, such as accelerometers and gyroscopes, to collect data. These sensors are practical and portable for use in any environment. However, the facial expression movement characteristics of users differ significantly, leading to noticeable variations in model performance across different users. Therefore, the system requires a small amount of data from the target user.
VI-A User-independent Model.
Through the experimental results, we found that the performance of the model after fine-tuning with user data is significantly better than that of the user-independent model. This is because the intensity, response, and sensitivity of facial movement data can vary significantly from person to person, leading to substantial differences between data domains. However, this may reduce usability, as new users must go through a registration phase before using the system. Our future research will explore generating large amounts of virtual IMU data using data from other modalities to train a more generalized user-independent model.
VI-B Effects of Body Movements.
In the system design, we considered the impact of initial head pose changes on the IMU signals. However, repetitive head bounces during activities such as walking can produce significant inertial signal noise [45], which can interfere with the normal operation of the system. Future research can explore the use of targeted filtering methods to minimize the noise impact introduced by such activities.
VII Conclusion

### v3

@@ 5.4.5 Latency and Power Consumption
In addition, we used a power monitoring device to measure the power consumption of IMUFace. Specifically, when the system is idle, that is, the microcontroller unit (MCU) is in sleep mode and does not transmit data via Bluetooth, the average power consumption of IMUFace is 50 mW (13.5 mA @ 3.70 V). When IMUFace actively collects IMU data and transmits it via Bluetooth, the total system power consumption is 58 mW (15.7 mA @ 3.70 V). This indicates that with a 100 mAh lithium polymer battery, IMUFace can continuously record data for up to 6.5 hours, meeting the needs of most applications. We examined the power consumption of recent studies on 3D facial reconstruction and summarized the overall power consumption of the system, including the use of sensing and communication power, in Table 3. Our analysis shows that IMUFace consumes the least power among all sensing technologies evaluated, with energy consumption 58.0% lower than that of the next most efficient technology, Bioface-3D [wu2021bioface].
Table 3: Comparison of System Power Consumption among Related Existing Studies
@@
NeckFace [chen2021neckface] Cameras 4 W
C-Face [chen2020c] Cameras >4 W
BioFace-3D [wu2021bioface] Biosensors 138 mW
EarIO [li2022eario] Acoustics 153.7 mW
EyeEcho [li2024eyeecho] Acoustics 167 mW
@@
6 Discussion
This study explores the use of conventional motion sensors, such as accelerometers and gyroscopes, for capturing facial expression data, emphasizing their practicality and portability for a wide range of applications. A key challenge identified is the variability in facial movements across different users, which impacts the model’s performance. To address this, incorporating a small amount of user-specific data has been shown to improve accuracy. This approach offers promising applications in fields like healthcare, gaming, and human-computer interaction by advancing real-time facial expression tracking and enabling innovative functionalities.
6.1 User-independent Model
Experimental results reveal that fine-tuning the model with user-specific data significantly enhances performance compared to a purely user-independent model. This improvement arises from the inherent variability in facial movement characteristics, such as intensity, responsiveness, and sensitivity, among individuals. These differences create unique data domains, posing challenges for generalized models. Although fine-tuning improves accuracy, it introduces a trade-off in usability, as new users must register and provide data for personalization.
To address this limitation, future work will focus on generating extensive virtual IMU datasets using multimodal inputs, such as video or depth sensors. By leveraging synthetic data, we aim to train a robust and generalized user-independent model capable of adapting to a broader user base with minimal personalization. Additionally, investigating techniques like domain adaptation and transfer learning could help bridge domain gaps, enhancing the model’s generalization capabilities. These strategies aim to maintain high performance while improving the system’s usability for diverse users.
6.2 Effects of Body Movements
While the system accounts for variations in initial head pose, dynamic body movements—such as repetitive head bounces during activities like walking or running—introduce significant noise into IMU signals [prakash2019stear]. This noise complicates the accurate detection of facial expressions, reducing the system’s reliability in real-world scenarios. To mitigate this issue, future research will explore advanced filtering techniques designed to isolate and remove noise caused by body movements while preserving subtle facial expression signals. Methods such as adaptive filtering, wavelet transforms, or machine-learning-based noise suppression may improve signal quality. Furthermore, integrating complementary sensors, such as magnetometers or pressure sensors, could provide additional context to distinguish between facial and body-induced movements. These enhancements aim to ensure reliable system performance in dynamic and mobile environments.
6.3 Practical Applications and Industry Implications
The proposed methodology demonstrates significant potential across various sectors, highlighting its practical value. In healthcare, it enables remote monitoring of patients with neurological disorders, facilitating the assessment of facial muscle activity and therapy progress. In gaming and virtual reality, it offers camera-free, real-time facial expression tracking, enhancing user privacy and reducing hardware complexity. For human-computer interaction, it supports more intuitive communication between users and devices. In the automotive industry, the system can monitor driver fatigue or emotional states, improving safety and user experience. These diverse applications underline the method’s versatility and its potential to address practical challenges across multiple disciplines.
7 Conclusion